OpenStudy (anonymous):

Three cards are drawn without replacement from a well-shuffled deck of 36 playing cards. What is the probability that the first card drawn is a heart?

OpenStudy (anonymous):

depends on how many hearts are in the 36 card deck most decks i know have 52 cards

OpenStudy (anonymous):

but if we assume that one quarter of them are hearts, then the probability the first card is a heart is \(\frac{1}{4}\)

OpenStudy (anonymous):

\[\frac{\heartsuit}{\heartsuit+\spadesuit+\clubsuit+\diamondsuit}=\frac{1}{4}\]

OpenStudy (anonymous):

I think it's supposed to look like the (A U B) problems, according to the book. I just have a problem with setting things up from a word problem.

OpenStudy (anonymous):

no, it is just the ratio of the number of hearts to the total number of cards that i sall